# FastEEC: Fast Evaluation of $N$-point Energy Correlators

### v1

@@ 5 Conclusions and outlook
In this letter, we focussed on the projected $N$-point energy correlator and proposed a method that provides a substantial speed up in the computation of higher point correlators. The underlying idea that we employ is that for correlations at a certain angular scale, radiation separated by much smaller distances can be treated as one, and radiation separated by larger distances can be treated as independent. We achieve this by reclustering with C/A or $k_{T}$, recursing over the tree, and using a dynamical subjet radius $r=\Delta R/\sqrt{f}$, with $\Delta R$ the separation between the two parents of the split under consideration. While C/A with fixed $f$ requires increasing the resolution to maintain accuracy for larger values of $N$, recustering with $k_{T}$ and using $f=k_{T,min}^{2}$ provides excellent accuracy for all higher-point correlators we studied. The gain in speed we obtain is one or more orders of magnitude, depending on the desired accuracy. We also utilize our method for the case where higher powers of transverse momenta are taken, which allows for a further speed up at the same relative accuracy.

### v2

@@ 5 Conclusions and outlook
In this Letter, we focused on the projected $N$-point energy correlator and proposed a method that provides a substantial speed up in the computation of higher point correlators. The underlying idea that we employ is that for correlations at a certain angular scale, radiation separated by much smaller distances can be treated as one, and radiation separated by larger distances can be treated as independent. We achieve this by reclustering with C/A or $k_{T}$, recursing over the tree, and using a dynamical subjet radius $r=\Delta R/\sqrt{f}$, with $\Delta R$ the separation between the two parents of the split under consideration. While C/A with fixed $f$ requires increasing the resolution to maintain accuracy for larger values of $N$, recustering with $k_{T}$ and using $f=k_{T,min}^{2}$ provides excellent accuracy for all higher-point correlators we studied. The gain in speed we obtain is one or more orders of magnitude, depending on the desired accuracy. We also utilize our method for the case where higher powers of transverse momenta are taken.